A geometric puzzle has a triangular puzzle piece with a perimeter of 44 centimeters. The length of the second side is double the first side. The length of the third side is 12 centimeters longer than the first side. Find the length of each side.
Answer by ikleyn(52778)

.

1-st side is x

2-nd side is 2x

3-rd side is (x+12).


The perimeter equation is


    x + 2x + x+12 = 44  centimeters.


Simplify and find x


       4x        = 44 - 12

       4x        = 32

        x = 8.


First side is  8 cm long;  second side is 16 cm long;  third side is 8+12 = 20 cm long.      ANSWER

Solved.